# =COLI-manual3.qry Cross-lingual Cross-temporal Summarization: Dataset, Models, Evaluation

### v1

@@ 6.1.1 The effect of different prompts and target language
A case study with ChatGPT-e2e vs. ChatGPT-pipeline. Though in Section 5 we observe better metric and human annotation results from ChatGPT e2e prompt, we notice that ChatGPT under this prompt is prone to output summaries in the wrong target language (invalid output). In such cases, we query the summary with the same prompt until a valid output is returned. We calculate the occurrences of invalid outputs after 2 rounds of queries and report them in Table 10, where we observe that using e2e prompts has a higher chance to yield invalid outputs. What is also worth noticing is when the output is expected in German, we observe more invalid cases among all three prompts which means ChatGPT tend to output English texts when our expected language is in German.
Model | Trg.lang | Invalid/Obs. |
@@ 6.1.3 ChatGPT knows it all?
Since our corpus-building process is highly dependent on Wikipedia (and popular stories presumably discussed on other places on the web), we want to know to what extent ChatGPT has an unfair advantage. We ask ChatGPT to output summaries based on the author, title and publication year. ChatGPT reports summaries to all requests in both language directions, though the output may contain irrelevant information besides the actual summarization.111111E.g. “The story emphasizes the power of kindness and the importance of inner beauty. It also highlights the value of perseverance and the rewards that can come from staying true to oneself.”
@@
However, despite the overall outstanding performance of ChatGPT summarization according to both humans, ChatGPT evaluation and automatic evaluation metrics, ChatGPT as a summarizer can also be prone to hallucination like other NLG systems Bang et al. (2023); Ji et al. (2023). For example, in the summary of the famous story Cinderella, we notice the golden slippers in the original document become glass slippers which is the best-known version worldwide but is actually incorrect based on the input text.
Adversarial attacks for ChatGPT
@@
Knowing that ChatGPT has specific prior knowledge of our CLCTS task, we take a step further and design three experiments inspired by the work from Chen and Eger (2023) where we gradually reduce the amount of specific prior knowledge in the input documents (from seen documents to pseudo-unseen documents to unseen documents). We aim to answer (i) does ChatGPT summarize truly based on the input text given its specific prior knowledge? (ii) Can
ChatGPT summarize against its specific prior knowledge? (iii) How well can ChatGPT summarize facing unseen documents?
@@
For question (i), we design sentence omission task. By dropping an increasing amount of sentences from the original documents, we control the information reduction of ChatGPT input for summarization. We randomly select 11 historical English documents from our CLCTS corpus containing 100-150 sentences and construct new documents by dropping sentences of different percentages. We then query summaries of the constructed documents through ChatGPT. To quantify the information change, we use the summary of the original documents (generated by ChatGPT) as the reference, and evaluate generated summaries from documents after omission using automatic evaluation metrics.121212We set the temperature to 0 since our goal is not to obtain the best summaries but to test ChatGPT summarization performance under gradual information loss. We average the resulting metric scores after scaling and use it as our similarity measure. The result is shown in Figure 4. We observe an overall downward trend of similarity score of both CTS and CLCTS which reflects a certain level of information loss accordingly. This indicates that ChatGPT summarizes according to the given information in general which contradicts our hypothesis that ChatGPT will be affected by its prior knowledge where the curve would be more flat.
@@
For question (ii), we first construct pseudo-unseen documents by entity swap, negation, and omission.131313For each attack, we construct 10 documents. For each document, we query the summaries 3 times at temperatures 0, 0.7 and 1. In total, we exclude summaries in the wrong target language and in the end, we have 257 valid generated summaries. Entity swap: We swap the entities from the original story with new entities. For example, in the story of Cinderella, we switch the hazel tree into cactus, or (to make it funny) we replace the prince with a watermelon (to ensure coherence, we also add background such as “They live in the Fruit Kingdom.”). Negation: We alter the plot completely to make a negated story. For example, in the story of Cinderella, the stepmother and stepsisters are mistreated by the rich man’s evil daughter and the prince eventually marries one of the stepsisters. Omission: For omission, we omit important plots. For example, Cinderella neither attends the ball nor marries the prince. We make sure each pseudo-unseen document only contains one kind of attack. We manually annotate the resulting output where we focus on if the attack is completed accurately or not (e.g., if during negation, Cinderella is mistreated again by the stepmother, then the task failed). We compute the accuracy and report it in Table 11.
@@
Omission | 0.74 | 0.63 |
Entity Swap | 0.80 | 0.71 |
Negation | 0.69 | 0.53 |
We observe from the results that (a) ChatGPT can better handle omission and entity swap than negation. This may be because compared to omission and entity swap, negation attacks alter the underlying logical connections of the story and thus requires more reasoning to complete this attack under the assumption that ChatGPT is affected by its explicit knowledge during summarization. (b) Higher accuracy for CLCTS than CTS task. For all three attacks, we observe a higher accuracy of completing the task for CTS. This is understandable since CLCTS also involves summarization cross-lingually which is more difficult.
To further investigate question (iii), we create 5 unseen documents from fiction chapters whose last updates are in year 2022 or later141414We include chatpters from 3 Chinese novels and 2 English novels.. For modern fiction chapters, we translate and rewrite the stories in historical languages and write reference summaries in modern languages. Both documents and summaries are proofread by a co-author of this work. Similar to Section 5, we evaluate the resulting summaries by three human annotators from four perspectives. The results are shown in Table 12. We notice that the overall rating for CTS and CLCTS ranges between 3 to 4 indicating that ChatGPT summarizes unseen documents at a mediocre level. This rating is similar to the results from our previous experiment using CLCTS corpus (we list the result of the same direction in Table 5).
ChatGPT can outperform finetuned models in terms of CLCTS tasks in our experiment. However, it does not magically solve all the issues as per our test, ChatGPT is better at omission and entity swap than negating against its prior knowledge and it generates summaries of mediocre quality from unseen source documents.
@@
unseen CTS hEn-En | 3.38 | 3.00 | 3.63 | 3.38 |
unseen CLCTS hEn-De | 3.40 | 3.17 | 3.78 | 3.17 |
our CLCTS hEn-De | 3.06 | 3.38 | 3.06 | 3.38 |
(a) Direction hDe-En
(b) Direction hEn-De
6.2 How well do different metrics perform at segment level?
Figure 5 shows Spearman’s ranking correlation between human annotation and evaluation metrics at segment level.
Decent correlation from BERTScore variants but worse for German For hDe-En, we observe a weak to moderate level of correlation with human annotation for BERTScore, and . However, the correlations for MENLI variants are negative. BARTScore correlates weakly in coherence and relevance. ROUGE-1 and ROUGE-L have low correlations with human annotation and the same is also true for MoverScore and DiscoScore. The situation is worse for hEn-De where we observe a very low level of correlation in general and less significant correlations from BERTScore variants according to the Student’s t-test but ROUGE-L correlates moderately in hEn-De in consistency and relevance. Overall, BERTScore variants correlate slightly better than other metrics with humans in hDe-En for coherence (0.49), consistency (0.51) and relevance (0.5), and in hEn-De for consistency (0.41) and relevance (0.26).
[…]
Additionally, we compute the correlation between evaluation metrics and human annotation among ChatGPT outputs only. Surprisingly, the resulting correlation is again moderate and even strong for BERTScore-P and DiscoScore (in consistency & relevance for hDe-En) though not significant according to a two-tailed Student’s t-test. However, this fact is still interesting since it may suggest that evaluation metrics work better when comparing outputs from similar training paradigms but worse when comparing between them (e.g., finetuned and prompt-based models).
@@ 7 Concluding remarks
In this work, we build the first CLCTS corpus and a historical translation dataset with additional information for future explorations. We study the characteristics of this corpus both alone and in comparison to other summarization corpora and examine the effectiveness of pipeline models, popular e2e transformer-based abstractive models leveraging intermediate task finetuning and ChatGPT for the CLCTS task. Even though these methods have been shown effective for single-task settings such as CLS or long document summarization, they fail to deliver good summaries in our combined task setting (cross-lingual, cross-temporal and long document). We show the potential of ChatGPT in CLCTS both as a summarizer and an evaluator coupled with different attacks to unveil its strength and weakness. Additionally, we find a moderate correlation of BERTScore variants with human annotation (slightly worse for German) and an overall improvement of correlation after excluding ChatGPT outputs at segment level.
Future work can address the limitations revealed by our experiments, including experiments with other models of great potential (e.g., unlimiformer Bertsch et al. (2023)), with other training frameworks (e.g., pertaining under multi-task learning) or with more historical data focusing on other languages (also more distant ones) and tasks (e.g., non-parallel historical corpora for pretraining and PAR3 cross-temporal dataset for machine translation Thai et al. (2022)). Furthermore, future work can focus on extending the size and diversity of our CLCTS datasets.
@@ 8 Acknowledgments
We thank all annotators for their hard work and time. Ran Zhang is financed by the BMBF project “Metrics4NLG”. Steffen Eger is financed by DFG Heisenberg grant EG 375/5–1.

### v3

@@ 6.3.1 The effect of different prompts and target language
A case study with ChatGPT-e2e vs. ChatGPT-pipeline. Though in Section 5 we observe better metric and human annotation results from ChatGPT e2e prompt, we notice that ChatGPT under this prompt is prone to output summaries in the wrong target language (invalid output). In such cases, we query the summary with the same prompt until a valid output is returned. We calculate the occurrences of invalid outputs after 2 rounds of queries and report them in Table 14, where we observe that using e2e prompts has a higher chance to yield invalid outputs. What is also worth noticing is when the output is expected in German, we observe more invalid cases among all three prompts which means that ChatGPT tends to output English texts when our expected language is in German.
Model | Tgt.lang | Invalid/Obs. |
@@ 6.3.3 ChatGPT knows it all?
Recently, the prevalence of LLMs has raised the concern of data contamination Balloccu et al. (2024); Sainz et al. (2023). Since our corpus-building process is highly dependent on Wikipedia (and popular stories presumably discussed on other places on the web), we want to know to what extent ChatGPT has an unfair advantage. We ask ChatGPT to output summaries based on the author, title, and publication year. ChatGPT reports summaries to all requests in both language directions, though the output may contain irrelevant information besides the actual summarization.212121For example, “The story emphasizes the power of kindness and the importance of inner beauty. It also highlights the value of perseverance and the rewards that can come from staying true to oneself.”
@@
However, despite the overall outstanding performance of ChatGPT summarization according to humans, ChatGPT evaluation, and automatic evaluation metrics, ChatGPT as a summarizer can also be prone to hallucinations like other NLG systems Bang et al. (2023); Ji et al. (2023). For example, in the summary of the famous story Cinderella, we notice the golden slippers in the original document become glass slippers which is the best-known version worldwide but is incorrect based on the input text.
@@
Adversarial attacks for ChatGPT Knowing that ChatGPT has specific prior knowledge of our CLCTS task, we go a step further and design three adversarial experiments inspired by the work from Chen and Eger (2023) where we gradually reduce the amount of specific prior knowledge in the input documents (from seen documents to pseudo-unseen documents to unseen documents). We aim to answer (1) does ChatGPT summarize truly based on the input text given its specific prior knowledge? (2) Can ChatGPT summarize against its specific prior knowledge? (3) How well can ChatGPT summarize facing unseen documents?
@@
For question (1), we design a sentence omission task. By dropping an increasing amount of sentences from the original documents, we control the information reduction of ChatGPT input for summarization. We randomly select 11 historical English documents from our CLCTS corpus containing 100-150 sentences and construct new documents by dropping sentences of varying percentages. We then use ChatGPT to summarize the new documents. To quantify the information change, we use the summary of the original documents (generated by ChatGPT) as the reference and evaluate generated summaries from documents after omission using automatic evaluation metrics.222222We set the temperature to 0 since our goal is not to obtain the best summaries but to test ChatGPT summarization performance under gradual information loss. We average the resulting metric scores after scaling and use it as our similarity measure. The result is shown in Figure 6. We observe an overall downward trend of similarity score of both CTS and CLCTS which reflects a certain level of information loss accordingly.
@@
For question (2), we first construct pseudo-unseen documents by entity swap, negation, and omission.232323For each attack, we construct 6 documents. We list examples of adversarial attacks in Section 9.9. For each document, we query the summaries 3 times at temperatures 0, 0.7, and 1 for both English and German. After excluding summaries in the wrong target language, we have 300 valid generated summaries. Entity swap: We swap multiple entities from the original story with new entities. For example, in the story of Cinderella, we switch the hazel tree into a cactus, and (to make it funny) we replace the prince with a watermelon (to ensure coherence, we also add background such as “They live in the Fruit Kingdom.”). Negation: We alter the important aspects of the plot to make a ‘negated story’. For example, in the story of Cinderella, the stepmother and stepsisters are mistreated by the rich man’s evil daughter and the prince eventually marries one of the stepsisters. Omission: For omission, we omit important plots. For example, Cinderella neither attends the ball nor marries the prince. We make sure each pseudo-unseen document only contains one kind of attack. We manually annotate the resulting output where we focus on whether the attack is completed accurately or not (e.g., if during negation, Cinderella is still mistreated by the stepmother — as in the original story — then the task is judged as a failure).242424We obtain the inter-annotator agreement between two annotators of 0.7 according to Cohen’s kappa and 0.87 according to percentage accuracy based on 15 samples. We list examples of failure cases in Section 9.10 (appendix). We compute the accuracy and report it in Table 15.
@@
Omission | 0.79 | 0.67 |
Entity Swap | 0.83 | 0.71 |
Negation | 0.76 | 0.53 |
We observe from the results that (1) ChatGPT can better handle omission and entity swap than negation. This may be because compared to omission and entity swap, negation attacks alter the underlying logical connections of the story and thus require more reasoning to complete this attack under the assumption that ChatGPT is affected by its explicit knowledge during summarization. (2) Higher accuracy for CTS than CLCTS task. For all three attacks, we observe a higher accuracy of completing the task for CTS. This is understandable since CLCTS also involves summarization cross-lingually which is more difficult.
To further investigate question (3), we create 15 unseen documents from fiction chapters whose last updates are in the year 2022 or later252525We include chapters from 5 Chinese novels and 10 English novels.. For modern fiction chapters, we translate and rewrite the stories in historical languages (by adding historical language elements such as morphological changes) and write reference summaries in modern languages. Both documents and summaries are proofread by a co-author of this work. Similar to Section 5, we evaluate the resulting summaries by three human annotators from four perspectives. The results are shown in Table 16. We notice that the overall rating for CTS and CLCTS ranges between 3 and 4 indicating that ChatGPT summarizes unseen documents at a mediocre level which is similar to the results from seen documents. This rating over all four dimensions for hEn-De is slightly worse than the results from our previous experiment using CLCTS corpus (we list the result of the same direction in row 3 in Table 16). We also query and annotate English summaries of unseen documents (CTS hEn-En) where we observe the overall performance is better than CLCTS hEn-De for unseen documents. This is expected since the CTS task does not involve cross-lingual summarization. However, we observe a decrease in summary quality of 0.75 points on average compared to the annotation results from CLCTS hDe-En task from our experiment. This is slightly surprising, since both target languages are in English, and especially CLCTS hDe-En has a more complex task setting. This may again indicate the benefits of the specific prior knowledge from ChatGPT.
@@
unseen CLCTS hEn-De | 3.24 | 2.79 | 3.06 | 2.85 |
(our CLCTS hEn-De | 3.31 | 2.80 | 3.38 | 2.98) |
unseen CTS hEn-En | 3.44 | 3.12 | 3.44 | 3.29 |
( our CLCTS hDe-En | 4.14 | 3.98 | 4.18 | 3.97) |
Overall, we argue that ChatGPT can clearly outperform finetuned models for CLCTS tasks in our experiments. However, it does not “magically” solve all the issues, as per our experiments: (1) it profits from its prior knowledge to a certain degree where it may also blend its prior knowledge with the observed data and ChatGPT is better at omission and entity swap than negating against its prior knowledge; (2) the benefit from prior knowledge slightly inflates its assessed quality where ChatGPT generates summaries of mediocre quality from unseen source documents with a slightly worse performance over all four evaluating dimensions compared to seen documents for hEn-De; (3) the performance is language dependent where compared to German outputs (hEn-De), English outputs (hEn-En) yield slightly better ratings.
@@ 7 Concluding remarks
In this work, we build the first CLCTS corpus and a historical translation dataset with additional information for future explorations. We study the characteristics of this corpus both alone and in comparison to other summarization corpora and examine the effectiveness of pipeline models, popular e2e transformer-based abstractive models leveraging intermediate task finetuning and ChatGPT for the CLCTS task. Even though these methods have been shown effective for single-task settings such as CLS or long document summarization, they fail to deliver good summaries in our combined task setting (cross-lingual, cross-temporal, and long document). We showcase the potential of GPT-3.5 in CLCTS as a zero-shot summarizer. It provides moderate to good quality outputs and seems very adept as a context-aware spelling normalization tool. However, it profits from its prior knowledge to a certain degree as tested by adversarially attacked and unseen source documents. Overall, we observe better performance for plot omission and entity swap than plot negation against its prior knowledge. Moreover, GPT-3.5 performs slightly worse for unseen source documents compared to seen documents.
As for evaluation, ChatGPT (both GPT-3.5 and GPT-4) as an evaluator can correlate with human evaluations at a moderate to good level but is prone to giving lower scores. Additionally, we find a moderate correlation of BERTScore with human annotation (slightly worse for German).
By regression analysis with BERTScore-F1, we quantify the effects of document features on model performance where we find positive impacts from documents with more recent publication years and document-summary embedding similarity while we observe a negative relation between model performance and the length of document. This helps explain model performance and affirms our hypotheses about the difficulty of the CLCTS task as discussed in the introduction (language change and divergence, etc.).
Future work can address the limitations revealed by our experiments. For example, researchers can explore other recent models of great potential, such as Unlimiformer under sufficient GPU memory Bertsch et al. (2024) and Longnet Ding et al. (2023). Other training frameworks such as pretraining under multi-task learning are also promising alternatives. Incorporating tasks such as non-parallel historical corpora for pretraining and the PAR3 cross-temporal dataset for machine translation Thai et al. (2022) may further provide benefits. Finally, future work can focus on extending the size and diversity of our CLCTS datasets, including more diverse language pairs.262626Our current work builds the CLCTS corpus for the high-resource language pair German and English. Main difficulties working with other languages include: (1) Even more Wikipedia summaries may be missing for other languages (German and English are two of the highest resource languages); (2) it may be more difficult to obtain historical sources of other languages; (3) human evaluation of other languages may be less accessible and thus even more costly; (4) automatic metrics including GPT4 may also be worse as we have observed worse performance for German than for English. While exploring more distant language pairs is thus immensely interesting, (5) our results further indicate that transfer would be even more difficult for all models involved.
@@ 8 Acknowledgments
We thank all annotators for their hard work and time. The NLLG group gratefully acknowledges support from the Federal Ministry of Education and Research (BMBF) via the grant “Metrics4NLG” and the German Research Foundation (DFG) via the Heisenberg grant EG375/5-1.